# October 20, 2025

### v1

@@ 5 Results
Figure 7: Extrapolated sensitivity to the optimised observables (from top left to bottom right) $\mathcal{H}_{2c}$, $\mathcal{Z}_{4}$, $S^{t}_{5}$, $Q_{4}^{\prime}$, $M_{2c}$, and $P_{5}^{t\prime}$ in comparison to the predictions of SM. The SM predictions are evaluated using flavio [47] with form-factors from Refs. [59, 60] and eos [61] where the form factors follow Refs. [62, 21, 63]. The variation of the accuracy for the different tagging-power setups within a given luminosity scenario is shown by the length between the endcap and the tip of the errorbar of the same colour. This affects $\mathcal{Z}_{4}$, $Q_{4}^{\prime}$, $S^{t}_{5}$, and $P_{5}^{t\prime}$ which are only measurable in the presence of flavour-tagging. In addition to the prediction in the SM, a alternative scenario is illustrated where the imaginary or real part of $\mathcal{C}_{9}$ is shifted by minus one. Small discrepancies at low-$q^{2}$ are observed and attributed to the massless lepton approximation used.
In order to assess how the additional mixing-observables improvement knowledge on the underlying short-distances physics mediating $Bs\!\rightarrow\phi$$\mu^{+}\mu^{-}$, we perform a global analysis of the measurements using the EFT framework. The analysis is similar to the global fits presented in Ref. [64]. The observables are parameterised depending on the Wilson coefficients and we profile the metric
@@
where $\vec{O}$ denotes the observable vector for the experiment and theory and $\underline{C}$ is the respective covariance matrix. We assume that there are no correlations between the experimental observables in the different $q^{2}$-regions and determine $\vec{O}_{\rm th.}(\mathcal{C}_{i})$ and $\underline{C}^{\rm th.}(\mathcal{C}_{i})$ using flavio.
Using this setup, we can compare the constraints of the Wilson-coefficients parameter space for the different luminosity milestones and for the different sets of observables present when including tagging or time-dependence in the PDF. The fits are performed using our fit results obtained in three $q^{2}$-regions from 0.10.980.1-0.98, 1.16.01.1-6.0, and $15.0-18.9\text{\,Ge\kern-1.00006ptV}^{2}\!/c^{4}$. Adding more $q^{2}$-regions will improve the constraints further. The constraints obtained when measuring the full time-dependent tagged decay rate for the different luminosity milestones are compared in Fig. 8 assuming a $6\,\%$ flavour-tagging power. Notably the the imaginary part of the Wilson coefficients can be better constrained than the real part when include the full-set of ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ observables.
@@
A comparison of the constraints obtained for the different fit-setups is shown in Fig. 9 for an integrated luminosity of $50\text{\,fb}^{-1}$ and a flavour-tagging power of $6\,\%$. The constraints on the imaginary part of the Wilson coefficients are improved when including the time-dependent observables and almost doubled when including the full set of observables accessible when including flavour-tagging. For the comparison of the time-dependent and time-independent setups, the $\mathcal{H}_{i}$ are removed from the list of observables that are fitted. The improvement in sensitivity from measuring the $\mathcal{H}_{i}$ in a single $q^{2}$-region is greater than that obtained when combining multiple $q^{2}$-regions. Figure 10 shows the constraints obtained when including the $\mathcal{H}_{i}$ and fitting only the data in the $q^{2}$-intervals from $0.1$ to $0.98\text{\,Ge\kern-1.00006ptV}^{2}\!/c^{4}$ and $1.1$ to $6.0\text{\,Ge\kern-1.00006ptV}^{2}\!/c^{4}$, respectively. An improved per-$q^{2}$ sensitivity to the Wilson coefficients is valuable because it enables more precise fits in the different $q^{2}$-regions, where the relative impact of non-local and non-factorisable SM contributions differ. The treatment of these non-factorisable contributions is under active scrutiny from the theory community, see for example Ref. [21], making it particularly valuable to study how the best fit-value of the Wilson Coefficients vary over $q^{2}$.
@@ 6 Conclusion and Discussion
We present the tagged time-dependent PDF for ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ and, by fitting this to realistic pseudo-experiments, we demonstrate for the first time that the complete set of angular observables associated with ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ can be extracted with $\mathcal{O}(10^{3})$ signal events under realistic assumptions for background levels and decay-time resolution expected at a hadronic machine. We show that this extraction can significantly improve constraints on the underlying Wilson coefficients, and that the sensitivity to their imaginary parts can be further enhanced through the optimized observables proposed in this work. Among these new observables, $Q_{4}^{\prime}$ and $M_{8}^{\prime}$ demonstrate a marked gain in sensitivity to short distance new physics effects, compared to their unoptimised counterparts. These observables should be measurable with a precision of around $0.3$ in the central-$q^{2}$ region by the end of the LHC era, assuming an effective tagging power between $4\text{--}6\,\%$. Of the tagging-accessible observables, $P_{5}^{t\prime}$ is particularly interesting, as its analogue in ${{B}^{0}}\!\rightarrow{{K}^{*0}}{\mu^{+}\mu^{-}}$ has exhibited significant tensions with the SM. We project that $P_{5}^{t\prime}$ can be measured with a precision of up to $0.1$ in the central $q^{2}$ region, again assuming an effective tagging power between $4\text{--}6\,\%$, by the end of the LHC era. Moreover, while the greatest sensitivity is achieved with a tagged analysis, we demonstrate that even an untagged time-dependent measurement, which is experimentally not as challenging as a tagged measurement, yields substantial improvements compared to the time-integrated analyses performed to date. This is especially true when examining the variation of the extracted Wilson coefficients across different $q^{2}$-regions, which can provide crucial insight into the origin of the shift in $\mathcal{R}\{C_{9}\}$ associated with the flavour anomalies.
@@ Acknowledgements
We express our gratitude to Martin Novoa-Brunet, Keri Vos, Peter Stangl, and Méril Reboud for the discussions about the implementation of the angular coefficients into flavio and eos. This project received support from the National Science Foundation grants CSSI-2411204 and PHY-2310073.
@@ Appendix B Time acceptance
The procedure is visualised in Figure 13, we find $A=1.05$, $B=13.3$, and $\delta_{t}=8.1\,\mathrm{ps}^{-1}$. Note that we neglect the uncertainties on the acceptance, as we are only interested in finding a function that captures the general behaviour of the time-acceptance. Our function is positive definite for all decay-times which are analysed, captures the kink in the points well and describes the plateau towards larger decay-times well.
Figure 13: Decay-time acceptance extracted from Ref. [52] with our best-fit overlayed. The function captures the behaviour of the acceptance well. The shaded area visualises the $68$-th percentiles of the acceptances obtained when randomly sampling the obtained parameters according to the uncertainties.
@@ Appendix F Sensitivity to the Optimised Observables
This appendix summarises the sensitivity to selected optimised observables and shows the predictions within the SM and when changing the vector-coupling $\mathcal{C}_{9}$. We choose to display observables which show sensitivity to changes in the vector coupling $\mathcal{C}_{9}$, which comprise $M_{4}^{(\prime)}$, $M_{8}^{(\prime)}$, $Q_{3}$, $Q_{4}^{(\prime)}$, $Q_{8}^{(\prime)}$, $Q_{9}$, $P_{2}^{t}$, $P_{4}^{t\prime}$, $P_{5}^{t\prime}$. To visualise the improved sensitivity to NP, also the un-optimised counterparts are shown. The $Q_{i}$ and $M_{i}$ are displayed in Fig. 22 and Fig. 23, respectively. The $P_{i}^{t}$ are shown in Fig. 24. The different luminosity scenarios are visualised using different colours and the different tagging power is indicated by the two errorbar sizes. Note that the tagging power influences variables which require tagging, such as the $Q_{i}$, whereas for example the $M_{i}$ are measurable without flavour-tagging. The definitions of the $M_{i}$ and $Q_{i}$ follow Eq. 21 and Eq. 26 and the $P_{i}^{t}$ are given in Eqs. 29 and following.
@@
Figure 24: Extrapolated sensitivities to $P_{2}^{t}$ (left), $P_{4}^{t\prime}$ (centre), $P_{5}^{t\prime}$ (right) and the corresponding un-optimised observables (bottom row). Overlayed are the predictions obtained within the SM and alternative scenarios where the vector coupling $\mathcal{C}_{9}$ is altered by minus one. Both $P_{2}^{t}$ and $P_{5}^{t\prime}$ are only measurable using flavour-tagging, hence their sensitivities depend on the tagging-power. The lower bound on the uncertainty, obtained using the better tagging-power, is given by the endcap of the errorbar, whereas the upper bound on the uncertainty is given by the upper end of the coloured errorbar.

### v2

@@ 5 Results
Figure 7: Extrapolated sensitivity to the optimised observables (from top left to bottom right) $\mathcal{H}_{2c}$, $\mathcal{Z}_{4}$, $S^{t}_{5}$, $Q_{4}^{\prime}$, $M_{2c}$, and $P_{5}^{t\prime}$ in comparison to the predictions of SM. The SM predictions are evaluated using flavio [49] with form-factors from Refs. [62, 63] and eos [64] where the form factors follow Refs. [65, 21, 66]. The variation of the accuracy for the different tagging-power setups within a given luminosity scenario is shown by the length between the end cap and the tip of the error bar of the same colour. This affects $\mathcal{Z}_{4}$, $Q_{4}^{\prime}$, $S^{t}_{5}$, and $P_{5}^{t\prime}$ which are only measurable in the presence of flavour-tagging. In addition to the prediction in the SM, an alternative scenario is illustrated where the imaginary or real part of $\mathcal{C}_{9}$ is shifted by minus one. Small discrepancies at low-$q^{2}$ are observed and attributed to the massless lepton approximation used.
In order to assess how the additional mixing-observables improve the knowledge on the underlying short-distances physics mediating ${{B}^{0}_{s}}\!\rightarrow\phi$$\mu^{+}\mu^{-}$, we perform a global analysis of the measurements using the EFT framework. The analysis is similar to the global fits presented in Ref. [67], which employs a weak effective Hamiltonian comprising a set of local dimension-six operators and effective coupling constants, the Wilson coefficients. In this work we focus on the four fermion $bs\mu\mu$ vector and axial-vector operators, as contributions to the other operators are better constrained by measurements using different decay modes. Additionally, whereas we focus on the left-chiral structure in this work, we expect similar constraining power for the right-chiral operators.
In order to evaluate the constraints on the parameter space, the observables are parameterised depending on the Wilson coefficients and we profile the metric
@@
where $\vec{O}$ denotes the observable vector for the experiment and theory and $\underline{C}$ is the respective covariance matrix. We assume that there are no correlations between the experimental observables in the different $q^{2}$-regions and determine $\vec{O}_{\rm th.}(\mathcal{C}_{i})$ and $\underline{C}^{\rm th.}(\mathcal{C}_{i})$ using flavio, evaluating the Wilson coefficients at the renormalisation scale $\mu=4.8\text{\,Ge\kern-1.00006ptV}$.
Using this setup, we can compare the constraints of the Wilson-coefficients parameter space for the different luminosity milestones and for the different sets of observables present when including tagging or time-dependence in the PDF. The fits are performed using our fit results obtained in three $q^{2}$-regions from 0.10.980.1-0.98, 1.16.01.1-6.0, and $15.0-18.9\text{\,Ge\kern-1.00006ptV}^{2}\!/c^{4}$. Adding more $q^{2}$-regions will improve the constraints further. The constraints obtained when measuring the full time-dependent tagged decay rate for the different luminosity milestones are compared in Fig. 8 assuming a $6\,\%$ flavour-tagging power. Notably the imaginary part of the Wilson coefficients can be better constrained than the real part when including the full-set of ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ observables.
@@
A comparison of the constraints obtained for the different fit-setups is shown in Fig. 9 for an integrated luminosity of $50\text{\,fb}^{-1}$ and a flavour-tagging power of $6\,\%$. The constraints on the imaginary part of the Wilson coefficients are improved when including the time-dependent observables and almost doubled when we include the full set of observables accessible if flavour-tagging is performed. For the comparison of the time-dependent and time-independent setups, the $\mathcal{H}_{i}$ are removed from the list of observables that are fitted. The improvement in sensitivity from measuring the $\mathcal{H}_{i}$ in a single $q^{2}$-region is greater than that obtained when combining multiple $q^{2}$-regions. Figure 10 shows the constraints obtained when including the $\mathcal{H}_{i}$ and fitting only the data in the $q^{2}$-intervals from $0.1$ to $0.98\text{\,Ge\kern-1.00006ptV}^{2}\!/c^{4}$ and $1.1$ to $6.0\text{\,Ge\kern-1.00006ptV}^{2}\!/c^{4}$, respectively. An improved per-$q^{2}$ sensitivity to the Wilson coefficients is valuable because it enables more precise fits in the different $q^{2}$-regions, where the relative impact of nonlocal and nonfactorisable SM contributions differ. The treatment of these nonfactorisable contributions is under active scrutiny from the theory community, see for example Ref. [21], making it particularly valuable to study how the best fit-value of the Wilson coefficients vary over $q^{2}$.
@@ 6 Conclusion and Discussion
We present the tagged time-dependent PDF for ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ and, by fitting this to realistic pseudoexperiments, we demonstrate for the first time that the complete set of angular observables associated with ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ can be extracted with $\mathcal{O}(10^{3})$ signal events under realistic assumptions for background levels and decay-time resolution expected at a hadronic machine. We show that this extraction can significantly improve constraints on the underlying Wilson coefficients, and that the sensitivity to their imaginary parts can be further enhanced through the optimized observables proposed in this work. Among these new observables, $Q_{4}^{\prime}$ and $M_{8}^{\prime}$ demonstrate a marked gain in sensitivity to short distance new physics effects, compared to their unoptimised counterparts. These observables should be measurable with a precision of around $0.3$ in the central-$q^{2}$ region by the end of the LHC era, assuming an effective tagging power between $4\text{--}6\,\%$. Of the tagging-accessible observables, $P_{5}^{t\prime}$ is particularly interesting, as its analogue in ${{B}^{0}}\!\rightarrow{{K}^{*0}}{\mu^{+}\mu^{-}}$ has exhibited significant tensions with the SM. We project that $P_{5}^{t\prime}$ can be measured with a precision of up to $0.1$ in the central $q^{2}$ region, again assuming an effective tagging power between $4\text{--}6\,\%$, by the end of the LHC era. Moreover, while the greatest sensitivity is achieved with a tagged analysis, we demonstrate that even an untagged time-dependent measurement, which is experimentally not as challenging as a tagged measurement, yields substantial improvements compared to the time-integrated analyses performed to date. This is especially true when examining the variation of the extracted Wilson coefficients across different $q^{2}$-regions, which can provide crucial insight into the origin of the shift in $\mathcal{R}\{C_{9}\}$ associated with the flavour anomalies.
@@ Acknowledgements
We express our gratitude to Martin Novoa-Brunet, Keri Vos, Peter Stangl, and Méril Reboud for the discussions about the implementation of the angular coefficients into flavio and eos. This project received support from the National Science Foundation Grants No. CSSI-2411204 and No. PHY-2310073.
Data Availability
The data that support the findings in this article are openly available [61], embargo periods may apply.
@@ Appendix B Time acceptance
The procedure is visualised in Figure 13, we find $A=1.05$, $B=13.3$, and $\delta_{t}=8.1\,\mathrm{ps}^{-1}$. Note that we neglect the uncertainties on the acceptance, as we are only interested in finding a function that captures the general behaviour of the time-acceptance. Our function is positive definite for all decay-times which are analysed, captures the kink in the points well and describes the plateau toward larger decay-times well.
Figure 13: Decay-time acceptance extracted from Ref. [54] with our best-fit overlayed. The function captures the behaviour of the acceptance well. The shaded area visualises the $68$th percentiles of the acceptances obtained when randomly sampling the obtained parameters according to the uncertainties.
@@ Appendix F Sensitivity to the Optimised Observables
This appendix summarises the sensitivity to selected optimised observables and shows the predictions within the SM and when changing the vector-coupling $\mathcal{C}_{9}$. We choose to display observables which show sensitivity to changes in the vector coupling $\mathcal{C}_{9}$, which comprise $M_{4}^{(\prime)}$, $M_{8}^{(\prime)}$, $Q_{3}$, $Q_{4}^{(\prime)}$, $Q_{8}^{(\prime)}$, $Q_{9}$, $P_{2}^{t}$, $P_{4}^{t\prime}$, $P_{5}^{t\prime}$. To visualise the improved sensitivity to NP, also the unoptimised counterparts are shown. The $Q_{i}$ and $M_{i}$ are displayed in Fig. 22 and Fig. 23, respectively. The $P_{i}^{t}$ are shown in Fig. 24. The different luminosity scenarios are visualised using different colours and the different tagging power is indicated by the two error bar sizes. Note that the tagging power influences variables which require tagging, such as the $Q_{i}$, whereas for example the $M_{i}$ are measurable without flavour-tagging. The definitions of the $M_{i}$ and $Q_{i}$ follow Eq. 21 and Eq. 26 and the $P_{i}^{t}$ are given in Eqs. 29 and following.
@@
Figure 24: Extrapolated sensitivities to $P_{2}^{t}$ (left), $P_{4}^{t\prime}$ (centre), $P_{5}^{t\prime}$ (right) and the corresponding unoptimised observables (bottom row). Overlayed are the predictions obtained within the SM and alternative scenarios where the vector coupling $\mathcal{C}_{9}$ is altered by minus one. Both $P_{2}^{t}$ and $P_{5}^{t\prime}$ are only measurable using flavour-tagging, hence their sensitivities depend on the tagging-power. The lower bound on the uncertainty, obtained using the better tagging-power, is given by the end cap of the error bar, whereas the upper bound on the uncertainty is given by the upper end of the coloured error bar.